# DeFi Liquidation Risk Modeling Using the Reflection Principle for Zero-Drift Brownian Motion

### v1

@@ Abstract
In this paper, we propose an analytical method to compute the collateral liquidation probability in decentralized finance (DeFi) stablecoin single-collateral lending. Our approach models the collateral exchange rate as a zero-drift geometric Brownian motion, converts it into a regular zero-drift Brownian motion, and employs the reflection principle to derive the liquidation probability. Unlike most existing methods that rely on computationally intensive simulations such as Monte Carlo, our formula provides a lightweight, exact solution. This advancement offers a more efficient alternative for risk assessment in DeFi platforms.
@@ 1 Introduction
In recent years, decentralized finance (DeFi) lending platforms have gained significant popularity as innovative alternatives to traditional financial services. These platforms enable users to borrow and lend digital assets in a transparent manner. One of the key concerns in DeFi lending is the risk of collateral liquidation, which can lead to substantial financial losses for lenders and borrowers alike. Collateral liquidation occurs when the collateral value falls below the liquidation threshold. Therefore, developing robust methods to assess and quantify the likelihood of liquidation events is essential to ensure the stability and security of DeFi ecosystems.
@@
Furthermore, existing methods for estimating liquidation probabilities often rely on computationally intensive techniques such as Monte Carlo simulations. Although these approaches can provide valuable information, they tend to be resource-intensive and time-consuming, limiting their practicality for real-time risk assessment and decision making. Developing analytical solutions that are accurate and computationally efficient remains an important goal in the field.
@@ 2 Deriving Analytical Solution
2.1 Modeling Collateral Exchange Rates as Zero-Drift Brownian Motion
Liquidation occurs when the value of the collateral drops below the liquidation threshold. Collateral value changes due to exchange rates, interest rates and the loan value, although exchange rate volatility diminishes the relevance of interest rates, allowing them to be disregarded in the context of mathematical modeling. We will model loans of stablecoins; therefore, we can assume that the loan value is constant and only model the collateral exchange rates. Also, we will model single-collateral loans.
Both DeFi and CeFi use relative change, not absolute change. This means that exchange rates are multiplied when they change instead of something being added to them or subtracted from them. This means that we can model the exchange rate as the following SDE:
@@ 2.1 Modeling Collateral Exchange Rates as Zero-Drift Brownian Motion
where $\sigma$ is the standard deviation or volatility, $S_{t}$ is the exchange rate and $W_{t}$ is a Wiener process. This is the equation of zero-drift geometric Brownian motion (GBM) [8] [7]. It is not convenient to work with this equation. So we divide it by $S_{t}$:
${\frac{dS_{t}}{S_{t}}}=\sigma dW_{t}$
Then we integrate both sides on $[0,t]$:
[…]
We get:
$ln\left(S_{t}\right)-ln\left(S_{0}\right)=\sigma W_{t}-\sigma W_{0}=\sigma W_{t}$
We can define $X_{t}$:
$X_{t}=ln\left(S_{t}\right)=\sigma W_{t}+ln\left(S_{0}\right)$
Now it’s a Brownian motion and a simple additive process:
$X_{t+\Delta t}=X_{t}+\sigma\sqrt{\Delta t}Z,\quad Z\sim\mathcal{N}\left(0,1\right)$
Or
$dX_{t}=\sigma dW_{t}$
@@ 2.2 Applying the Reflection Principle
Since we disregard interest rates, we can define the liquidation threshold and the exchange rate at which collateral value reaches the liquidation threshold to be constant. Therefore, we can use the reflection principle.
The reflection principle tells us about the probability of a Wiener process crossing some value in some time from now [6]. Here it is:
@@
Where $\Phi$ is the cumulative distribution function of standard normal distribution. Now we have a complete analytical solution to our problem.
@@ 3 Simulating Collateral Exchange Rates
Monte Carlo simulation [3] [4] [5] can be used to validate the theoretical formula derived earlier. To simulate collateral exchange rate paths, we model the evolution of the price as a stochastic process with Gaussian noise using a small timestep $dt$. There are two common ways to discretize the GBM for simulation:
@@ 4 Simulation Results and Validation
Figure 1: Probability calculated by the reflection principle and probability calculated by Monte Carlo.
Days Theory GBM Euler-Maruyama Abs. Dif. GBM Abs. Dif. Euler-Maruyama
3 0.003288 0.00184 0.00231 0.001448 0.000978
7 0.054313 0.03556 0.04241 0.018753 0.011903
14 0.173605 0.13239 0.15020 0.041215 0.023405
30 0.352608 0.30196 0.33991 0.050648 0.012698
60 0.510997 0.46529 0.51849 0.045707 0.007493
90 0.591495 0.55254 0.61697 0.038955 0.025475
120 0.642095 0.60637 0.67139 0.035725 0.029295
Table 1: Comparison of liquidation probabilities and their absolute differences from the theoretical values
Figure 1 shows the liquidation probability within the given time in days calculated using the formula derived earlier, the probabilities calculated by Monte Carlo with one hundred thousand simulations for each method and the absolute difference between theory and simulation for each method. The starting exchange rate is 1500 and the exchange rate at which the liquidation happens is 1212. Standard deviation or annualized volatility $\sigma=0.8$. From the results we see that the formula derived closely aligns with the simulations. From the figure we see that the difference doesn’t rise with the increase in time.
@@ 6 Conclusion
In this paper, we derived and validated a closed-form analytical formula for computing liquidation probability in DeFi lending protocols. By modeling the evolution of cryptocurrency prices as a zero-drift geometric Brownian motion and applying the reflection principle from probability theory, we were able to express the likelihood of liquidation in a compact, non-simulation-based form. This method avoids the computational overhead of traditional Monte Carlo simulations while retaining strong accuracy across various time horizons. Through a thorough numerical comparison, we showed that the theoretical estimates match closely with empirical results from both Euler–Maruyama and exact GBM-based simulations, with error remaining stable and small.

### v2

# DeFi Liquidation Risk Modeling Using Geometric Brownian Motion

@@ Abstract
In this paper, we propose an analytical method to compute the collateral liquidation probability in decentralized finance (DeFi) stablecoin single-collateral lending. Our approach models the collateral exchange rate as a zero-drift geometric Brownian motion, and derives the probability of it crossing the liquidation threshold. Unlike most existing methods that rely on computationally intensive simulations such as Monte Carlo, our formula provides a lightweight, exact solution. This advancement offers a more efficient alternative for risk assessment in DeFi platforms.
@@ 1 Introduction
In recent years, decentralized finance (DeFi) lending platforms have gained significant popularity as innovative alternatives to traditional financial services. These platforms enable users to borrow and lend digital assets in a transparent manner. One of the key concerns in DeFi lending is the risk of collateral liquidation, which can cause substantial financial losses for lenders and borrowers alike. Collateral liquidation occurs when the collateral value falls below the liquidation threshold. Therefore, developing robust methods to assess and quantify the likelihood of liquidation events is essential to ensure the stability and security of DeFi ecosystems.
@@
Furthermore, existing methods for estimating liquidation probabilities often rely on computationally intensive techniques such as Monte Carlo simulations. Although these approaches can provide valuable information, they tend to be resource intensive and time consuming, limiting their practicality for real-time risk assessment and decision making. Developing analytical solutions that are accurate and computationally efficient remains an important goal in the field.
@@
2 Exploratory Data Analysis
After collecting and organizing information on the daily changes in the opening and closing exchange rates of a selected cryptocurrency, we conducted an initial exploratory analysis to better understand its behavior over time. One of the first visualizations we produced was a histogram representing the distribution of daily rate changes (Figure 1).
From the shape of the histogram, it is evident that the data approximately follows a bell-shaped curve, resembling the normal distribution. This suggests that most of the daily changes cluster around the mean, with fewer extreme values occurring at the tails.
400-400200-200020020040040060060001122334455667788991010111112121313141415151616171718181919202021212222232324242525262627272828292930303131323233333434353536363737383839394040Exchange Rate Change Close - Open Distribution for Period: 29.03.2024 - 29.03.2025
Figure 1: Histogram of daily change in Ethereum exchange rate
@@ 3 Deriving Analytical Solution
3.1 Liquidation Criteria
Liquidation occurs when the value of the collateral drops below the liquidation threshold. Collateral value changes due to exchange rates, interest rates and the loan value, although exchange rate volatility diminishes the relevance of interest rates, allowing them to be disregarded in the context of mathematical modeling. We will model loans of stablecoins; therefore, we can assume that the loan value is constant and only model the collateral exchange rates. In addition, we will model single-collateral loans.
3.2 Modeling Collateral Exchange Rates as Zero-Drift Geometric Brownian Motion
A classic way to quantitatively model financial assets is Geometric Brownian Motion (GBM) [9] [8]. It also fits the collected data. We will use it to model our exchange rates. It has the following stochastic differential equation (SDE):
@@ 3.2 Modeling Collateral Exchange Rates as Zero-Drift Geometric Brownian Motion
where $\sigma$ is the standard deviation or volatility, $S_{t}$ is the exchange rate, and $W_{t}$ is a Wiener process. We can rewrite it as:
$dS_{t}=x\left(t,S_{t}\right)dt+y\left(t,S_{t}\right)dW_{t}$
where $x\left(t,S_{t}\right)=0$ and $y\left(t,S_{t}\right)=\sigma S_{t}$. We can define $f\left(Y_{t}\right)=ln\left(Y_{t}\right)$ and now we can apply Itô lemma to get:
$df=\left(\frac{\partial f}{\partial t}+x\frac{\partial f}{\partial S_{t}}+\frac{y^{2}}{2}\frac{\partial^{2}f}{\partial S_{t}^{2}}\right)dt+y\frac{\partial f}{\partial S_{t}}dW_{t}$
Since $f\left(Y_{t}\right)=ln\left(Y_{t}\right)$, $\frac{\partial f}{\partial t}=0$, $\frac{\partial f}{\partial S_{t}}=\frac{1}{S_{t}}$ and $\frac{\partial^{2}f}{\partial S_{t}^{2}}=-\frac{1}{S_{t}^{2}}$ After simplifying we get:
$d\left(ln\left(S_{t}\right)\right)=df=\sigma dW_{t}-\frac{\sigma^{2}}{2}dt$
We can integrate from 0 to $t$ over $dt$ and get:
$ln\left(Y_{t}\right)-ln\left(Y_{0}\right)=\sigma W_{t}-\frac{\sigma^{2}}{2}t$
Define $X_{t}=ln\left(Y_{t}\right)$:
$X_{t}=X_{0}+\sigma W_{t}-\frac{\sigma^{2}}{2}t$
which is a drifting Brownian motion. For small periods of time or small volatility, drift can be assumed to be zero, and approximation can be achieved via the reflection principle. Otherwise, we will have to use the inverse Gaussian distribution.
@@ 3.3 Applying the Reflection Principle
Since we disregard interest rates, we can define the liquidation threshold and the exchange rate at which the collateral value reaches the liquidation threshold to be constant. Therefore, we can use the reflection principle.
Currently, we will disregard the drift $\frac{\sigma^{2}}{2}t$ in the GBM equation to approximate a solution as we said earlier. Later, we will see for what time periods this approximation works.
The reflection principle tells us about the probability that a Wiener process will cross some value in some time [7]. Here it is:
@@
Where $\Phi$ is the cumulative distribution function (CDF) of standard normal distribution. Now we have a complete analytical solution to our problem. We have derived the closed-form solution for our approximation.
3.4 Applying the Inverse Gaussian Distribution
The inverse Gaussian distribution $IG\left(\mu_{0},\lambda\right)$ can be thought as the distribution of the first hitting time of a threshold $a>0$ for Brownian motion[3]:
$\begin{array}[]{l}X_{0}=0\\ X_{t}=\sigma W_{t}+\mu t\;,\;\mu>0\end{array}$
We can ensure $X_{0}=0$ by shifting the coordinates to fit $X_{0}=0$. We get $a_{1}=X_{\text{liquidation}}-X_{0}=ln\left(S_{\text{liquidation}}\right)-ln\left(S_{0}\right)=ln\left(\frac{S_{\text{Liquidation}}}{S_{0}}\right)<0$. We know $\mu_{1}=-\frac{\sigma^{2}}{2}<0$. These are negative but have the same sign. Now, flip the coordinate system, and we get $a_{2}=-a_{1}=ln\left(\frac{S_{0}}{S_{\text{Liquidation}}}\right)>0$ and $\mu_{2}=-\mu_{1}=\frac{\sigma^{2}}{2}$. The first passage time probability density function (PDF) of such process has the following formula:
$f\left(t_{1}\right)=\frac{a_{2}}{\sigma\sqrt{2\pi t^{3}}}exp\left(-\frac{\left(a_{2}-\mu_{2}t\right)^{2}}{2\sigma^{2}t}\right)$
for
$t_{1}=inf\left\{\tau>0\;|\;X_{\tau}=a_{2}\right\}\sim IG\left(\frac{a_{2}}{\mu_{2}},\;\left(\frac{a_{2}}{\sigma}\right)^{2}\right)$
We want to predict the probability of it crossing the threshold in time $t$ from the start. Since the PDF of the inverse Gaussian distribution is 0 for negative inputs, we only need to compute the CDF for $t$. CDF for $IG\left(\mu_{0},\lambda\right)$ at the value $x$ is:
$F\left(x\right)=\Phi\left(\sqrt{\frac{\lambda}{x}}\left(\frac{x}{\mu_{0}}-1\right)\right)+exp\left(\frac{2\lambda}{\mu_{0}}\right)\Phi\left(-\sqrt{\frac{\lambda}{x}}\left(\frac{x}{\mu_{0}}+1\right)\right)$
Replacing $x$ with $t$, $\mu_{0}$ with $\frac{a_{2}}{\mu_{2}}$ and $\lambda$ with $\left(\frac{a_{2}}{\sigma}\right)^{2}$ results in the probability of liquidation.
@@ 4 Simulating Collateral Exchange Rates
Monte Carlo simulation [4] [5] [6] can be used to validate the theoretical formula derived earlier. To simulate collateral exchange rate paths, we model the evolution of the price as a stochastic process using a small time step $dt$. There are two common ways to discretize the GBM for simulation:
@@
    $S_{t+dt}=S_{t}\exp\left(\sigma\sqrt{dt}Z-\frac{\sigma^{2}}{2}dt\right)$
  • Zero-drift GBM solution: Since we have neglected drift for the reflection principle approximation. We will also simulate it:
@@ 5 Simulation Results and Validation
Row Annualized Volatility Starting Exchange Rate Liquidation Exchange Rate $\frac{S_{0}}{S_{\text{liquidation}}}$
1 0.8 1500 1200 1.25
2 0.8 2500 1200 2.083333
3 1.8 1500 1200 1.25
4 1.8 2500 1200 2.083333
5 2.8 1500 1200 1.25
6 2.8 2500 1200 2.083333
Table 1: Simulation values
@@
Figure 2: Simulation results and comparison of methods. Probability of liquidation over the time frame in days. Simulation values can be found in table 1. 10000 paths were simulated for each method with a timestep of 1 day.
Figure 2 shows the simulation results. As expected, the formula derived from the reflection principle works well for low volatility and smaller time frames, but performs poorly under high volatility. The difference between the GBM and Euler-Maruyama simulations and the formula also grows with time. However, the difference between the formula and the no-drift GBM simulation stays consistent. We also see that the formula derived from the inverse Gaussian distribution generally shows more accurate results.
@@ 7 Conclusion
In this paper, we derived and validated a closed-form analytical formula for computing liquidation probability in DeFi lending protocols. By modeling the evolution of cryptocurrency prices as a zero-drift geometric Brownian motion and applying the reflection principle and inverse Gaussian distribution from probability theory, we were able to express the likelihood of liquidation in a compact, non-simulation-based form. This method avoids the computational overhead of traditional Monte Carlo simulations while retaining strong accuracy across various time horizons. Through a thorough numerical comparison, we showed that the theoretical estimates match closely with empirical results from both Euler–Maruyama and exact GBM-based simulations, with error remaining stable and small.